OpenStudy (anonymous):

I graphed y=3x+1 on a number graph and now I need to write an equation for the line parallel to the line on that graph that passes through (2,0) can anyone help me do this step by step?

OpenStudy (aum):

y = 3x + 1 What is the slope of this line?

OpenStudy (anonymous):

(1,0)

OpenStudy (anonymous):

if thats correct

OpenStudy (aum):

That is not the slope. The general equation of a line in the slope-intercept form is y = mx + b. 'm' is the slope and 'b' is the y-intercept.

OpenStudy (anonymous):

i wan't told to solve the problem just to graph it, so I typed the equation into mathway and I was given the points to plot it on

OpenStudy (anonymous):

so no slope

OpenStudy (aum):

But this question asks for the equation of a line parallel to the line y = 3x + 1 and passing through the point (2,0). To find the equation we now need to first find the slope of the line y = 3x + 1.

OpenStudy (anonymous):

i'm going to send a picture to help if i may

OpenStudy (aum):

That is part A. The graph has to be steeper than that. Doesn't look right.

OpenStudy (anonymous):

part B is right below it! that's what this question was about, and I believe that is right due to google and mathway telling me thats the way it should look but it could be wrong

OpenStudy (aum):

y = 3x + 1 represents a straight line. To draw a straight line all you need are ANY two points on the line. You don't need mathway or Google to do that. Just find two points on the line. y = 3x + 1 put x = 0: y = 3*0 + 1 = 1 (0, 1) is one point. put x = 1: y = 3*1 + 1 = 4 (1, 4) is another point. Plot those two points on the graph paper. Then using a ruler draw a line passing through those two points.

OpenStudy (anonymous):

alright i just need to plot the points one minute

OpenStudy (aum):

I will come back after for you post the pic of the corrected graph.

OpenStudy (anonymous):

excuse the bad handwriting

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Correct.

OpenStudy (aum):

For part B) they want an equation of a line parallel to y = 3x + 1 and passing through the point (2,0). Parallel lines have the SAME slope. What is the slope of the line y = 3x + 1. Compare this equation to the general form y = mx + b where 'm' is the slope. we can see 3 is the slope of the line y = 3x + 1. Therefore, the slope of the line we want is 3 so that it will be parallel. Put m = 3 in y = mx + b y = 3x + b This line has to pass through (2,0) That is, when x = 2, y = 0 0 = 3*2 + b 0 = 6 + b b = -6 y = 3x - 6 is the equation of the line parallel to the line y = 3x + 1 and passes through the point (2,0).

OpenStudy (anonymous):

i apologize if i dont respond right away the website is acting up but keep going please!

OpenStudy (anonymous):

may i show you part c quick in case i keep crashing, unless you dont have time.

OpenStudy (aum):

It is done. They just want the equation and it is solved above. They don't want graph for this one.

OpenStudy (aum):

go ahead.

OpenStudy (aum):

C) Two lines are perpendicular if the product of their slopes is -1. Slope of the line y = 3x + 1 is 3. Let y = mx + b be the slope of the perpendicular line. The slope of this line is m. Since the two lines are perpendicular, the product of their slopes, 3 * m = -1 Therefore, m = -1/3 The perpendicular line is: y = -1/3x + b It passes through (0, -4) when x = 0, y = -4 -4 = -1/3 * 0 + b b = -4 y = -1/3x - 4 is the equation of the perpendicular line.

OpenStudy (anonymous):

this is starting to make sense now! do you think you have time for the bonus question as well?

OpenStudy (aum):

I will do one more.

OpenStudy (anonymous):

wait I think you made a mistake in the last question

OpenStudy (anonymous):

at the last part did you mean to put when x = 0, y = -4 -4 = -1/3 * 0 + b b = -4 y = -1/3x - 4 is the equation of the perpendicular line.

OpenStudy (anonymous):

there are two -4 and bs in a row

OpenStudy (anonymous):

it's confusing me a bit :(

OpenStudy (anonymous):

@aum no rush just making sure you know i'm still here

OpenStudy (anonymous):

wait i got it!

OpenStudy (anonymous):

heres the bonus question

OpenStudy (anonymous):

oh god thats small

OpenStudy (aum):

can't see.

OpenStudy (anonymous):

it says: the standard form of a linear equation is Ax+By=C. On another page, transform this into the slope-intercept form using A,B,C,x, and y

OpenStudy (aum):

Ax + By = C (Standard form of a linear equation). To put it in slope-intercept form, we need to put it in the form y = mx + b. Ax + By = C Solve for y. Isolate y: Add -Ax to both sides: By = -Ax + C divide by B: \[y = -\frac ABx + \frac CB \] is the slope-intercept form.

OpenStudy (anonymous):

awesome : D thank you so much for your help i made you best response and fanned you dont worry

OpenStudy (aum):

you are welcome.